OpenStudy (anonymous):

Use synthetic division and the given factor to comletely factor each polynomial function y=x^3-3x^2-10x+24; (x-2). I divided it and got 1x^2+1x-12,Remainder 48.. Is this right and how do i factor it with the remainder?

OpenStudy (lgbasallote):

what did you divide it with again?

OpenStudy (anonymous):

so i divided (x^3-3x^2-10x+24) by (x-2)

OpenStudy (lgbasallote):

i think you went wrong...there shouldn't be a remainder

OpenStudy (anonymous):

Hmmmmm I thought so ... let me try again

OpenStudy (lgbasallote):

i have to go now so good luck

OpenStudy (anonymous):

Thanks anyway!

jimthompson5910 (jim_thompson5910):

Use synthetic division to get what you see attached

jimthompson5910 (jim_thompson5910):

The first three numbers at the bottom represent the coefficients of the quotient, so the quotient is x^2 - x - 12 The last number is the remainder, which is 0. So (x^3-3x^2-10x+24)/(x-2) = x^2 - x - 12 and you can rewrite the equation to get x^3-3x^2-10x+24 = (x-2) (x^2 - x - 12)

OpenStudy (anonymous):

I was doing that but i see i got my positive and negative signs tumbled up. Now i would just factor 1x^2-1x-12?

jimthompson5910 (jim_thompson5910):

yes exactly

jimthompson5910 (jim_thompson5910):

what do you get when you factor that

OpenStudy (anonymous):

okay hold on just a moment let me try and do this (:

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

would it be (x+3)(x-4)?

jimthompson5910 (jim_thompson5910):

yes it is, (x+3)(x-4) x(x-4) + 3(x-4) x^2 - 4x + 3x - 12 x^2 - x - 12 So follow this backwards to factor x^2 - x - 12 into (x+3)(x-4)

jimthompson5910 (jim_thompson5910):

So x^3-3x^2-10x+24 completely factors to (x-2)(x+3)(x-4)

OpenStudy (anonymous):

Awesome!!! Okay i did that right! Thank you so much for your help (: (:

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

:) :)